OpenStudy (anonymous):

Determine whether the graph of y = x2 − 6x + 3 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (-6, 3) Maximum; (-6, 3) Minimum; (3, -6) Maximum; (3, -6)

OpenStudy (jdoe0001):

it's a parabola,and it has a SQUARED "x" variable, meaning is moving vertically the "a" component is positive, meaning is opening upwards how do you find the vertex point?, well \(\bf \left(-\cfrac{b}{2a}, c-\cfrac{b^2}{4a}\right)\)

OpenStudy (anonymous):

right ? \[\left(\begin{matrix}-6 \\ 2x\end{matrix}\right)\]\[3 -\frac{ 6^2 }{ 4x }\]

OpenStudy (jdoe0001):

hmm \(\Large\begin{matrix} y = &x^2& − 6x& + 3\\ y = &1x^2& − 6x& + 3\\ &a\ \ \ &b\ \ \ &c \end{matrix}\)

OpenStudy (jdoe0001):

so \(\bf \left(-\cfrac{(-6)}{2(1)}, 3-\cfrac{(-6)^2}{4(1)}\right)\)

OpenStudy (anonymous):

ugh idk :(

OpenStudy (jdoe0001):

well, what's \(\bf -\cfrac{(-6)}{2(1)}\)

OpenStudy (jdoe0001):

or \(\bf \cfrac{-1(-6)}{2(1)}\) if you prefer :)

OpenStudy (anonymous):

2

OpenStudy (jdoe0001):

6/2 = 2? so what's 4/2 then?

OpenStudy (anonymous):

no 3

OpenStudy (jdoe0001):

so x = 3 what about \(\bf \cfrac{(-6)^2}{4(1)}\)

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

-6

OpenStudy (jdoe0001):

\(\bf \cfrac{(-6)^2}{4(1)} \implies \cfrac{-6 \times -6}{4 \times 1}\)

OpenStudy (anonymous):

36 over 4

OpenStudy (jdoe0001):

yes, so if you simplify 36/4?

OpenStudy (anonymous):

i got 9

OpenStudy (jdoe0001):

ahhh 9, so \(\bf 3-\cfrac{(-6)^2}{4(1)} \implies 3-(9)\)

OpenStudy (jdoe0001):

so the vertex is at x = 3, y = -6 (3, -6)

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

minimum or maximum

OpenStudy (jdoe0001):

it's a parabola,and it has a SQUARED "x" variable, meaning is moving vertically the "a" component is positive, meaning is opening upwards

OpenStudy (jdoe0001):

so what do you think? minimum or maximum?

OpenStudy (jdoe0001):

minimum is how LOW a function gets in an interval maximum is how HIGH it gets in an interval so you can think of them this way a minimum is a burrow a maximum is a hump

OpenStudy (anonymous):

minimum

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

:) thank you

OpenStudy (jdoe0001):

np